OpenStudy (anonymous):

Use a calculator to evaluate cot pi/5

OpenStudy (anonymous):

\[\cot \left(\frac{\pi }{5}\right)=\sqrt{1+\frac{2}{\sqrt{5}}}=1.37638 \]

OpenStudy (idealist10):

1.37638